OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (anonymous):

equation of the line when u have slope n point is y-y1=slope(x-x1)

OpenStudy (anonymous):

where (x1 y1)=(2,-4) use them n proceed

OpenStudy (anonymous):

y-(-4)=slope(x-2)

OpenStudy (anonymous):

put val of slope also n find equation

OpenStudy (anonymous):

Oh right, forgot about the slope value. y-(-4)=3/5(x-2)

OpenStudy (anonymous):

3x-5y=(-4)?

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

Thanks!